# Conversational Query Reformulation with the Guidance of Retrieved Documents

### v3

@@ Abstract
Conversational search seeks to retrieve relevant passages for the given questions in conversational question answering. Conversational Query Reformulation (CQR) improves conversational search by refining the original queries into de-contextualized forms to resolve the issues in the original queries, such as omissions and coreferences. Previous CQR methods focus on imitating human written queries which may not always yield meaningful search results for the retriever. In this paper, we introduce GuideCQR, a framework that refines queries for CQR by leveraging key information from the initially retrieved documents. Specifically, GuideCQR extracts keywords and generates expected answers from the retrieved documents, then unifies them with the queries after filtering to add useful information that enhances the search process. Experimental results demonstrate that our proposed method achieves state-of-the-art performance across multiple datasets, outperforming previous CQR methods. Additionally, we show that GuideCQR can get additional performance gains in conversational search using various types of queries, even for queries written by humans. ***The code and pre-trained model will be publicly released upon the acceptance of the paper.
@@ 1 Introduction
In this paper, we propose GuideCQR, an effective CQR framework designed to generate retrieval-friendly conversational queries by leveraging the guidance of initially retrieved documents from the base query. GuideCQR first obtains the initially retrieved documents through a retrieval process using the baseline query reformulated by LLMs. We then perform a re-ranking process to refine the order of the retrieved documents based on their similarity to the query. Through this re-ranking process, we carefully select a small number of documents from the initially retrieved ones to extract signals that are more likely to contain the gold passage. Next, we extract the keywords and generate the expected answer from the re-ranked guided documents to obtain components for creating a retriever-friendly query. We then independently filter both the keywords and the expected answers based on their similarity to the original query and previous utterances to eliminate redundant information. Finally, we construct the final query set by concatenating the filtered keywords and expected answer with the baseline query.
Experimental results show that GuideCQR achieves state-of-the-art across several CQR datasets compared to the baseline systems. We also demonstrate the robustness of the GuideCQR framework in enhancing the retrievability of queries, making them more retriever-friendly. Furthermore, we validate the effectiveness of our method by analyzing the overlap between the augmented keywords and the relevant documents for the given question, confirming the alignment between the augmented queries and document content. Additionally, we show that the proposed method achieves notable improvements across various query sets, underscoring its adaptability.
@@ 2.1 Conversational Search
To address such challenges in conversational search, two main methods that enable conversational search are conversational dense retrieval and conversational query reformulation. Conversational dense retrieval focuses on enhancing the representation of the current query by incorporating historical context achieved through training dense retrievers Qu et al. (2020). On the other hand, conversational query reformulation transforms the conversational search into a traditional ad-hoc search by converting the entire search session into a single standalone query Elgohary et al. (2019).
@@ 2.2 Conversational Query Reformulation
To address these challenges, recent studies have introduced methods like CONQRR Wu et al. (2022), which directly optimizes query rewriting for retrieval through Self-Critical Sequence Training Rennie et al. (2017), and ConvGQRMo et al. (2023a), which improves retrieval performance by combining query rewriting with query expansion. Additionally, IterCQR Jang et al. (2024) iteratively trains the CQR model by using information retrieval (IR) signals as rewards. The integration of large language models (LLMs) in CQR has also shown promise in recent work Mao et al. (2023).
@@ 3 GuideCQR
We propose GuideCQR, a novel framework designed to reformulate conversational queries by utilizing guidance from initially retrieved documents. Figure 2 provides an overview of the proposed query reformulation process. GuideCQR consists of three stages: We first retrieve an initial set of guided documents using a query reformulated by LLM. This step retrieves documents that are likely to contain gold passages to guide the query reformulation process (Sec 3.1). Next, we extract keywords and generate expected answers from the guided documents, creating components that contribute to making the query more retriever-friendly (Sec 3.2). Finally, we apply a filtering process that evaluates both keywords and answers based on their similarity scores relative to the baseline query and dialogue history. We then unify and concatenate these components back to the baseline query (Sec 3.3).
@@ 3.1.1 Initial Documents Retrieval
where represents an operation that resolves omissions or coreferences using OpenAI gpt3.5-turbo-16k and denotes the raw question in the dataset. denotes the dialogue history of , especially consisting of previous dialogues’ queries, rewritten queries and responses by human annotators. Based on both the and , we generate the . Using this , we obtain the initial documents composed of 2,000 documents for each sample ID, which are crucial for creating our final query set.
@@ 3.1.2 Re-ranking Documents
To further improve the quality of the guided set of documents, we implement a re-ranking process to generate initially retrieved documents. By reordering the initial documents using a different retrieval model, we aim to capture better documents by reducing biases that may arise from relying on a single retriever. Specifically, we employ Sentence-Transformer Reimers and Gurevych (2019) to re-rank the top 2,000 documents for each sample ID, selecting the final guided set by choosing the top 10 documents based on their similarity scores to the query.
@@ 3.2 Generating Retriever-friendly Query
Since many retrieved documents may contain influential words or signals that significantly impact subsequent retrieval stages, identifying key elements from these documents is crucial for constructing more effective queries. Based on the guided documents obtained from the previous step, we generate retriever-friendly queries with respect to two aspects: Augmenting Keywords, Expected Answer generation
@@ 3.2.1 Augmenting Keywords
Keywords play a critical role in forming retriever-friendly queries, as they capture the most relevant and significant terms within the document. To obtain these keywords, we utilize the guided documents.
We extract keywords by leveraging BERT Devlin et al. (2019) using the keyword extraction method as in KeyBERT Grootendorst (2020). We use BERT to compute embeddings, creating a representation for the entire document. Subsequently, we extract embeddings for N-gram words/phrases in the query. By computing cosine similarity, we determine which words/phrases are most similar to the document. We choose the keywords with the highest similarity.
In line with this principle, we augment the base queries with keywords through two key aspects. The first aspect concerns determining the number of documents to extract keywords from, which defines the level of guidance we aim to achieve. For each document, we extract keywords of a specified span length. The second aspect focuses on the span length of the keywords, which refers to the number of tokens, or the number of keywords, selected per document. For example, when augmenting top-2 span-3 keywords, we extract three keywords from each of the top two documents, resulting in a total of six keywords. Using this approach, we augment a total keyword list composed of keywords from the top- documents and span length :
where denotes unit keyword, is the number of documents and is keyword span length. For example, keywords such as "early", "throat", "cancer", "high", and "cure" can be crucial for the document as shown in Figure 2.
@@ 3.2.2 Expected Answer Generation
Expected answers for the given query can also serve as valuable signals to the query, providing direct insights into the user’s intent and thereby improving the relevance and accuracy of search results. We find that when generating the expected answer, rather than instructing language models to generate potential answers to the query, guided documents may contain gold answers to the query and can play a crucial role as a pool for finding the answers to reformulate query.
Hence, we also utilize the guided documents in generating expected answers as additional information to the query:
where represents answer list and denotes a unit answer extracted from a single document. To generate these answers, both the query and relevant context are necessary. We use the query as the baseline query and the context as the guided documents. For each document, we generate a single expected answer, such as "throat cancer has a high cure rate," as illustrated in Figure 2. We then aggregate these answers into a single sentence, treating the entire set as a unified response. Since each span contributes meaningfully to the final output, we avoid analyzing individual tokens individually. Consequently, this process produces k distinct answers, with one derived from each of the top-k documents.
@@ 3.3 Filtering and Unify
We observe that redundant elements, such as the keyword ’rate’ as shown in Figure 2, can emerge from both the augmented keywords and the generated answers. We explain this is because GuideCQR might augment the keywords or answers generated from irrelevant documents. So, irrelevant signals from irrelevant documents can have a negative impact when creating retriever-friendly queries.
To address this, we introduce an additional filtering stage to more effectively remove irrelevant keywords and answers from the reformulated query. This filtering process is guided by a metric we define as FilterScore, which leverages both and , calculated using cosine similarity, as follows:
@@ 4.1 Datasets and Metrics
We utilize three CQR benchmark datasets TREC CAsT-19 (Dalton et al., 2020), TREC CAsT-20 (Dalton et al., 2021) and QReCC dataset Anantha et al. (2021) for our work. CAsT-19 and CAsT-20 consist of 50 and 25 conversations, respectively. Unlike CAsT-19, CAsT-20 is more realistic and complex because its queries are based on information needs drawn from commercial search logs, and they can reference prior system responses. CAsT series datasets assign a relevance score ranging from 1 to 4 to each passage, indicating the degree of relevance to the query, with documents scoring a 4 being considered gold passages. For the QReCC dataset, each query is paired with a single gold passage different from CAsT dataset. QReCC dataset includes a training set and a test set, and we sample 2K conversations from the training set to create a development set, following the previous work Kim and Kim (2022).
@@
4.2 Baselines
We compare GuideCQR with the following CQR methods: (1) Transformer++ (Vakulenko et al., 2021a): A GPT-2 based CQR model fine-tuned on CANARD Elgohary et al. (2019) dataset. (2) CQE-Sparse (Lin et al., 2021b): A weakly-supervised method to select important tokens only from the context via contextualized query embeddings. (3) QuReTeC (Voskarides et al., 2020b): A weakly-supervised method to train a sequence tagger to decide whether each term contained in a historical context should be added to the current query. (4) T5QR (Lin et al., 2020): A conversational query rewriter based on T5, trained using human-generated rewrites. (5) ConvGQR (Mo et al., 2023a): A query reformulation framework that combines query rewriting with generative query expansion. (6) CONVINV (Cheng et al., 2024): Framework that transforms conversational session embeddings into interpretable text using Vec2Text. (7) CHIQ Mo et al. (2024): A two-step method that leverages the capabilities of LLMs to resolve ambiguities in the conversation history before query rewriting. (8) IterCQR Jang et al. (2024): Improves a CQR framework through iterative training method based on the similarity between the passage and the query. (9) LLM4CS: Query rewriting based on LLM and various prompting methods. We build GuideCQR upon this prompted QR method using (REW + Maxprob) and evaluate using self-consistency.
@@ 4.5 Analysis
Performance among Amount of Initial Guided documents
We evaluate the impact and robustness of each step in the GuideCQR setup. Initially, we adjust the number of guided documents to observe the proper quantity and present the results in Table 3. Our findings demonstrate that increasing the number of guided documents consistently enhances performance. However, retrieving an excessive number of documents leads to longer inference times. Therefore, we limit the number of initial documents to 2,000 for our main experiment.
| # of initial guided documents | MRR | NDCG@3 | Avg |
| 10 | 75.7 | 50.5 | 63.1 |
| 100 | 78.2 | 51.4 | 64.8 |
| 1,000 | 79.5 | 52.1 | 65.8 |
| 2,000 | 81.0 | 52.2 | 66.6 |
@@
As presented in Table 4, we examine the impact of varying the number of documents used in augmenting keywords and generating expected answers. For instance, when extracting keywords and answers from the top two documents, we generate one keyword-answer pair per document, resulting in two pairs in total. Our findings indicate that incorporating a larger number of documents generally provides additional information, which can improve performance. However, using too many documents may cause a decline in overall performance similar to the amount of initial guided documents.
@@
Performance among Keyword Span Length
| CAsT-19 | CAsT-20 | |||
| MRR | NDCG@3 | MRR | NDCG@3 | |
| 72.2 | 45.2 | 53.4 | 37.2 | |
| 5 span | 76.9 | 49.9 | 57.7 | 41.5 |
| 10 span | 78.8 | 52.0 | 53.7 | 39.8 |
| 15 span | 81.0 | 52.2 | 54.1 | 40.0 |
| 20 span | 80.1 | 52.0 | 54.6 | 39.8 |
@@
where represents the number of unique matched keywords found in the guided documents that have query relevance score of 4 and represents the total number of unique augmented keywords. Note that the document that has a relevance score of 4 is regarded as the most relevant document with user queries in CAsT datasets. Matching keywords in these documents can confirm that augmenting keywords plays a crucial role in generating retriever-friendly queries. As shown in Table 6, a longer keyword span results in a higher precision score, highlighting the significance of keyword augmentation. However, an excessive number of documents may negatively impact the overall score.
| Span Length | Precision |
| 5 | 0.6684 |
| 10 | 0.7018 |
| 15 | 0.7024 |
| 20 | 0.7045 |
@@
GuideCQR is a highly effective framework as it operates as a query expansion method, making it adaptable for use with other methods or queries simultaneously. To demonstrate its effectiveness, we conduct experiments using both human-rewritten queries and from the CAsT-19 dataset. The refers to the original query without any reformulation applied. As shown in Table 7, both and human-rewritten queries show performance improvements. This demonstrates that our framework can be applied to any query or CQR method to achieve enhanced performance.
| MRR | NDCG@3 | |
| 41.2 | 24.3 | |
| + GuideCQR | 47.0(+5.8) | 28.8(+4.5) |
| Human Rewrite | 74.0 | 46.1 |
| + GuideCQR | 82.8(+8.8) | 53.0(+6.9) |

### v4

@@ Abstract
Conversational search seeks to retrieve relevant passages for the given questions in conversational question answering. Conversational Query Reformulation (CQR) improves conversational search by refining the original queries into de-contextualized forms to resolve the issues in the original queries, such as omissions and coreferences. Previous CQR methods focus on imitating human written queries which may not always yield meaningful search results for the retriever. In this paper, we introduce GuideCQR, a framework that refines queries for CQR by leveraging key information from the initially retrieved documents. Specifically, GuideCQR extracts keywords and generates expected answers from the retrieved documents, then unifies them with the queries after filtering to add useful information that enhances the search process. Experimental results demonstrate that our proposed method achieves state-of-the-art performance across multiple datasets, outperforming previous CQR methods. Additionally, we show that GuideCQR can get additional performance gains in conversational search using various types of queries, even for queries written by humans. 111The code and pre-trained model will be publicly released upon the acceptance of the paper.
@@ 1 Introduction
In this paper, we propose GuideCQR, an effective CQR framework designed to generate retrieval-friendly conversational queries by leveraging the guidance of initially retrieved documents from the base query. GuideCQR first obtains the initially retrieved documents through a retrieval process using the baseline query reformulated by LLM. We then perform a re-ranking process to refine the order of the retrieved documents based on their similarity to the query. Through this re-ranking process, we carefully select a small number of documents from the initially retrieved ones to extract signals that are more likely to contain the gold passage. Next, we extract the keywords and generate the expected answer from the re-ranked guided documents to obtain components for creating a retriever-friendly query. We then independently filter both the keywords and the expected answers based on their similarity to the original query and previous utterances to eliminate redundant information. Finally, we construct the final query set by concatenating the filtered keywords and expected answers with the baseline query.
Experimental results show that GuideCQR achieves state-of-the-art across several CQR datasets compared to the baseline systems. We also demonstrate the robustness of the GuideCQR framework in enhancing the retrievability of queries, making them more retriever-friendly. Furthermore, we validate the effectiveness of our method by analyzing the overlap between the augmented keywords and the relevant documents for the given question. Additionally, we show that the proposed method achieves notable improvements across various query sets, underscoring its adaptability.
@@ 2.1 Conversational Search
To address such challenges in conversational search, two main methods that enable conversational search are Conversational Dense Retrieval (CDR) and CQR. CDR focuses on enhancing the representation of the current query by incorporating historical context achieved through training dense retrievers Qu et al. (2020). On the other hand, CQR transforms the conversational search into a traditional ad-hoc search by converting the entire search session into a single standalone query Elgohary et al. (2019).
@@ 2.2 Conversational Query Reformulation
To address these challenges, recent studies have introduced methods like CONQRR Wu et al. (2022), which directly optimizes query rewriting for retrieval through Self-Critical Sequence Training Rennie et al. (2017), and ConvGQRMo et al. (2023a), which improves retrieval performance by combining query rewriting with query expansion. Additionally, IterCQR Jang et al. (2024) iteratively trains the CQR model by using information retrieval (IR) signals as rewards. The integration of LLMs in CQR has also shown promise in recent work Mao et al. (2023).
@@ 3 GuideCQR
We propose GuideCQR, a novel framework designed to reformulate conversational queries by utilizing guidance from initially retrieved documents. Figure 2 provides an overview of the proposed query reformulation process. GuideCQR consists of three stages: We first retrieve an initial set of guided documents using a query reformulated by LLM. This step retrieves documents that are likely to contain gold passages to guide the query reformulation process (Sec 3.1). Next, we extract keywords and generate expected answers from the guided documents, creating components that contribute to making the query more retriever-friendly (Sec 3.2). Finally, we apply a filtering process that evaluates both keywords and answers based on their similarity scores relative to the baseline query and dialogue history. We then unify and concatenate these components with the baseline query to construct the final query for post retrieval (Sec 3.3).
@@ 3.1.1 Initial Documents Retrieval
where represents an operation that resolves omissions or coreferences using OpenAI gpt3.5-turbo-16k and denotes the raw question in the dataset, without any reformulation applied. denotes the dialogue history of , especially consisting of previous dialogues’ queries, rewritten queries and responses by human annotators. Based on both the and , we generate the . Using this , we obtain the initial documents composed of 2,000 documents for each question. They are crucial for creating our final query set and strong foundation for identifying meaningful signals and guiding the retriever towards the most relevant passages.
@@ 3.1.2 Re-ranking Documents
To further improve the quality of the guided set of documents, we implement a re-ranking process to generate initially retrieved documents. By reordering the initial documents using a different retrieval model, we aim to capture better documents by reducing biases that may arise from relying on a single retriever. Specifically, we employ Sentence-Transformer Reimers and Gurevych (2019) to re-rank the top 2,000 documents for each question, selecting the final guided set by choosing the top 10 documents based on their similarity scores to the query.
@@ 3.2 Generating Retriever-friendly Query
Since many retrieved documents may contain influential words or signals that significantly impact subsequent retrieval stages, identifying key elements from these documents is crucial for constructing more effective queries. Based on the guided documents obtained from the previous step, we generate retriever-friendly queries by incorporating extra information to the query from two approaches: Augmenting Keywords, Expected Answer generation
@@ 3.2.1 Augmenting Keywords
We find that keywords from the initial documents play a critical role in forming retriever-friendly queries, as they capture the most relevant and significant terms within the document. For example, keywords such as "early", "throat", "cancer", "high", and "cure" can be beneficial if they are augmented to the search query as shown in Figure 2.
We leverage KeyBERT Grootendorst (2020) to extract keywords from the re-ranked documents. The process begins by using BERT Devlin et al. (2019) to compute embeddings of documents, which create a representation for the entire document. Embeddings for N-gram words or phrases within the document are then extracted. By calculating the cosine similarity between these words/phrases and the document representation, the method identifies which words/phrases are most similar to the document.
In line with this principle, we enhance the base queries by augmenting keywords through two hyperparameters. The first one involves determining the number of documents to extract keywords from, which establishes the level of guidance we intend to provide. For each guided document, we extract keywords of a specified span length. The second aspect pertains to the span length, which indicates the number of tokens or keywords selected per document. For instance, when augmenting with top-2 span-3 keywords, we extract three keywords from each of the top two documents, yielding a total of six keywords.
Consequently, we augment a total keyword list composed of keywords from the top- documents and span length :
where denotes unit keyword, is the number of documents and is keyword span length.
@@ 3.2.2 Expected Answer Generation
Guided documents often include gold answers to the query, serving as a valuable resource for efficiently reformulating the query. Based on this idea, we use guided documents as context to generate expected answers to enhance the query. As illustrated in Figure 2, although it’s not obtained from the gold passage, expected answer is a concise and informative response that can potentially address the user’s query since they can provide direct insights into the user’s intent, thereby improving the relevance and accuracy of search results. Specifically, we generate these expected answers as follows:
where represents answer list and denotes a unit answer extracted from a single document. To generate these answers, both the query and relevant context are necessary. We use the query as the baseline query and the context as the guided documents. We generate a single expected answer for each document. Consequently, this process produces k distinct answers, with one derived from each of the top-k documents.
@@ 3.3 Filtering and Unify
We observe that redundant elements, such as the keyword ’rate’ as shown in Figure 2, can emerge from both the augmented keywords and the generated answers. We find that this is because GuideCQR might augment keywords or answers derived from irrelevant documents. Hence, irrelevant signals from irrelevant documents can have a negative impact when creating retriever-friendly queries.
To address this, we introduce an additional filtering stage to more effectively remove irrelevant keywords and answers from the reformulated query. We guide this filtering process through the metric FilterScore, which leverages both and , calculated using cosine similarity, as follows:
@@ 4.1 Datasets and Metrics
We utilize three CQR benchmark datasets TREC CAsT-19 (Dalton et al., 2020), TREC CAsT-20 (Dalton et al., 2021) and QReCC dataset Anantha et al. (2021) for our work. CAsT-19 and CAsT-20 consist of 50 and 25 conversations, respectively. Both CAsT datasets share the same document collection and provide passage-level relevance judgments, as well as human rewrites for each turn. Unlike CAsT-19, CAsT-20 is more realistic and complex because its queries are based on information needs drawn from commercial search logs, and they can reference prior system responses. CAsT series datasets assign a relevance score ranging from 1 to 4 to each passage, indicating the degree of relevance to the query, with documents scoring a 4 being considered gold passages. For the QReCC dataset, each query is paired with a single gold passage different from CAsT dataset. QReCC dataset includes a training set and a test set, and we sample 2K conversations from the training set to create a development set, following the previous work Kim and Kim (2022).
@@ 4.2 Implementation Details
4.3 Baselines
We compare GuideCQR with the following CQR methods: (1) Transformer++ (Vakulenko et al., 2021a): A GPT-2 based CQR model fine-tuned on CANARD Elgohary et al. (2019) dataset. (2) CQE-Sparse (Lin et al., 2021b): A weakly-supervised method to select important tokens only from the context via contextualized query embeddings. (3) QuReTeC (Voskarides et al., 2020b): A weakly-supervised method to train a sequence tagger to decide whether each term contained in a historical context should be added to the current query. (4) T5QR (Lin et al., 2020): A conversational query rewriter based on T5, trained using human-generated rewrites. (5) ConvGQR (Mo et al., 2023a): A query reformulation framework that combines query rewriting with generative query expansion. (6) CONVINV (Cheng et al., 2024): Framework that transforms conversational session embeddings into interpretable text using Vec2Text. (7) CHIQ Mo et al. (2024): A two-step method that leverages the capabilities of LLMs to resolve ambiguities in the conversation history before query rewriting. (8) IterCQR Jang et al. (2024): CQR framework through iterative refinement based on the similarity between the passage and the query. (9) LLM4CS: Query rewriting based on LLM and various prompting methods.
@@ 4.5 Analysis
CAsT dataset includes multiple relevant passages, each with a relevance score. We focus on CAsT due to these unique label features.
@@
As shown in Table 3, we verify the impact of varying the number of documents used in augmenting keywords and generating expected answers. Our findings indicate that incorporating a larger number of documents generally provides additional information, which can improve performance. However, using too many documents may cause a decline in overall performance similar to the amount of initial guided documents.
@@
Performance among Amount of Initial Guided documents
| # of initial guided documents | MRR | NDCG@3 | Avg |
| 10 | 75.7 | 50.5 | 63.1 |
| 100 | 78.2 | 51.4 | 64.8 |
| 1,000 | 79.5 | 52.1 | 65.8 |
| 2,000 | 81.0 | 52.2 | 66.6 |
We evaluate the impact and robustness of each step in the GuideCQR setup. Initially, we adjust the number of guided documents to observe the proper quantity and present the results in Table 4. Our findings demonstrate that increasing the number of guided documents consistently enhances performance. However, retrieving an excessive number of documents leads to longer inference times. Hence, considering the computation cost, we decide to use 2000 documents which is the point on the dev set where there is no additional performance improvement in at least one metric. We also use the same criteria for keyword span length.
GuideCQR MRR NDCG@3 R@10 MRR NDCG@3 R@10 RawQuery 41.2 24.3 5.8 47.0(+5.8) 28.8(+4.5) 6.6(+0.8) Human Rewrite 74.0 46.1 12.9 82.8(+8.8) 53.0(+6.9) 14.4(+1.5) GPT3.5-turbo-16k 72.2 45.2 12.2 81.8(+9.6) 53.7(+8.5) 13.9(+1.7) Qwen2.5-7B-Instruct 62.9 39.3 10.2 65.0(+2.1) 42.3(+3.0) 11.7(+1.5) Mistral-7B-Instruct-v0.3 69.1 42.8 11.6 69.8(+0.7) 46.0(+3.2) 12.0(+0.4) EXAONE-3.5-7.8B-Instruct 66.4 41.5 11.1 72.6(+6.2) 46.7(+5.2) 12.5(+1.4)
@@
where represents the number of unique matched keywords found in the guided documents that have query relevance score of and represents the total number of unique augmented keywords. Note that the document that has a relevance score of 1 or is regarded as the most relevant document with user queries in CAsT datasets. Matching keywords in these documents can confirm that augmenting keywords plays a crucial role in generating retriever-friendly queries. As shown in Table 6, the precision score for relevant passages is higher than that for irrelevant passages, indicating that augmenting keywords play a crucial role in retrieving relevant documents.
| Precision | ||
| Irrelevant | Passage (0) | 57.21 |
| Relevant | Passage (1) | 67.39 |
| Passage (2) | 71.36 | |
| Passage (3) | 72.26 | |
| Passage (4) | 70.45 |
@@
GuideCQR is a highly effective framework as it operates as a query expansion method, making it adaptable for use with other methods or queries simultaneously. To demonstrate its effectiveness, we conduct experiments using human-rewritten queries, and with different generator models from the CAsT-19 dataset. We create using GPT-3.5-turbo-16k to derive the final performance of GuideCQR. Furthermore, we conduct experiments with EXAONE-3.5-7.8B-Instruct Research (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), and Qwen2.5-7B-Instruct Team (2024). As shown in Table 5, , human-rewritten queries and with different generators show performance improvements. This demonstrates that our framework can be applied to any query or CQR method to achieve enhanced performance.